OpenStudy (anonymous):

How many grams of CaSO4 are needed to prepare an 875 mL solution with a concentration of 2.0 M ?

OpenStudy (anonymous):

do you need the number or the solution?

OpenStudy (anonymous):

Molar mass of CaSO4 = 136.1406 g/mol \[875 ml \times \frac{ 2 mole }{ 1 liter } = 1750 mmol\] of CaSO4 now \[1750 mmol \times \frac{ 136 gram }{ 1 mol } = 238 grams\] of CaSO4